### v3

@@ Abstract
This note proposes a new reformulation of complex semidefinite programs (SDPs) as real SDPs. As an application, we present an economical reformulation of complex SDP relaxations of complex polynomial optimization problems as real SDPs and derive some further reductions by exploiting inner structure of the complex SDP relaxations. Various numerical examples demonstrate that our new reformulation runs significantly faster than the usual popular reformulation.
@@ 1 Introduction
Complex semidefinite programs (SDPs) arise from a diverse set of areas, such as combinatorial optimization [9], optimal power flow [10, 12], quantum information theory [2, 4, 17], signal processing [11, 14]. In particular, they appear as convex relaxations of complex polynomial optimization problems (CPOPs), giving rise to the complex moment-Hermitian-sum-of-squares (moment-HSOS) hierarchy [10, 15, 16]. However, most modern SDP solvers deal with only real SDPs [^1] As far as the author knows, SeDuMi [13], Sdolab [7], and Hypatia [6] are the only solvers that can handle complex SDPs directly.. In order to handle complex SDPs via real SDP solvers, it is mandatory to reformulate complex SDPs as equivalent real SDPs. A popular way [^2] See for instance the online modeling cookbook of the commercial SDP solver MOSEK: https://docs.mosek.com/modeling-cookbook/sdo.html. to do so is to use the equivalent condition
@@
for an Hermitian matrix variable $H=H_{R}+H_{I}\hbox{\bf{i}}\in{\mathbb{C}}^{n\times n}$ with $H_{R}$ and $H_{I}$ being its real and imaginary parts respectively. Note that the right-hand-side constraint in (1) entails certain structure and to feed it to an SDP solver, we need to impose extra affine constraints to the positive semidefinite (PSD) constraint $Y\succeq 0$:
@@
This conversion is quite simple but could be inefficient when $n$ is large. In this note, inspired by Lagrange duality, we propose a new reformulation of complex SDPs as real SDPs. The benefit of this new reformulation is that there is no need to add extra affine constraints and hence it owns lower complexity. In the same manner, we can obtain a new reformulation of complex SDP relaxations of CPOPs as real SDPs. Moreover, by exploiting inner structure of the complex SDP relaxations, we are able to remove a bunch of redundant affine constraints, which leads to an even more economical real reformulation of the complex SDP relaxations. Various numerical experiments (including randomly generated CPOPs and the alternating current optimal power flow (AC-OPF) problem) confirm our theoretical expectation and demonstrate that the new reformulation is more efficient than the usual popular one. Actually, our implementation of the new reformulation with MOSEK [1] also runs much faster than the implementation of the original complex formulation with Hypatia [8], probably because the SDP solvers based on real numbers are more mature and robust.
[…]
@@
$\langle u,v\rangle_{{\mathbb{R}}}={\mathcal{R}}(u^{{\mathrm{H}}}v)={\mathcal{R}}(u)^{\intercal}{\mathcal{R}}(v)+{\mathcal{I}}(u)^{\intercal}{\mathcal{I}}(v),\quad u,v\in{\mathbb{C}}^{m}.$
For $u,v\in{\mathbb{R}}^{m}$, $\langle u,v\rangle\coloneqq u^{\intercal}v$, where $u^{\intercal}$ stands for the transpose of $u$. For each $A\in{\mathbb{C}}^{n\times n}$, we associate it with an Hermitian matrix ${\mathcal{H}}(A)\coloneqq\frac{1}{2}(A+A^{{\mathrm{H}}})$. One can check that ${\mathcal{R}}(\langle A,H\rangle)=\langle{\mathcal{H}}(A),H\rangle$ for any $H\in{\mathrm{H}}^{n}$.
@@ 2 The real reformulations of complex SDPs
($-{\mathbb{C}}$) $\begin{cases}\sup\limits_{H\in{\mathrm{H}}^{n}}&\langle C,H\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}(H)=b,\\ &H\succeq 0,\end{cases}$
where $C\in{\mathrm{H}}^{n},b\in{\mathbb{C}}^{m}$. In order to convert ($-{\mathbb{C}}$) to a real SDP, we define two real linear operators ${\mathscr{A}}_{R},{\mathscr{A}}_{I}\colon{\mathbb{R}}^{n\times n}\rightarrow{\mathbb{R}}^{m}$ associated to ${\mathscr{A}}$ by
@@
respectively. Moreover, assume $H=H_{R}+H_{I}\hbox{\bf{i}},C=C_{R}+C_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$ with $H_{R},H_{I},C_{R},C_{I}\in{\mathbb{R}}^{n\times n}$, $b_{R},b_{I}\in{\mathbb{R}}^{m}$. We can now convert ($-{\mathbb{C}}$) to a real SDP:
($-{\mathbb{R}}$) $\begin{cases}\sup\limits_{Y\in{\mathbf{S}}^{2n}}&\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(H_{R})+{\mathscr{A}}_{I}(H_{I})=b_{R},\\ &{\mathscr{A}}_{R}(H_{I})-{\mathscr{A}}_{I}(H_{R})=b_{I},\\ &Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\succeq 0.\end{cases}$
As mentioned in the introduction, to feed the PSD constraint in ($-{\mathbb{R}}$) to an SDP solver, we need to include also the extra $n(n+1)$ affine constraints listed in (2), which could be inefficient in practice. Below we show that by taking a dual point of view, we can actually get rid of this issue.
Before formulating the dual problem of ($-{\mathbb{C}}$), we explicitly give the adjoint operator of ${\mathscr{A}}$.
@@
Then by the convex duality theory, the dual problem of ($-{\mathbb{C}}$) reads as
($-{\mathbb{C}}$) $\begin{cases}\inf\limits_{y\in{\mathbb{C}}^{m}}&\langle b,y\rangle_{{\mathbb{R}}}\\ \,\,\,\rm{s.t.}&{\mathscr{A}}^{*}(y)\succeq C.\end{cases}$
@@
Thus, we can convert ($-{\mathbb{C}}$) to a real SDP by using the equivalent condition (1):
($-{\mathbb{R}}$) $\begin{cases}\inf\limits_{y_{R},y_{I}\in{\mathbb{R}}^{m}}&b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}\\ \,\,\,\,\,\,\,\,\rm{s.t.}&\begin{bmatrix}U-C_{R}&-V+C_{I}\\ V-C_{I}&U-C_{R}\end{bmatrix}\succeq 0.\end{cases}$
Let $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]\in{\mathbf{S}}^{2n}$ be the dual PSD variable of ($-{\mathbb{R}}$) with $X_{1},X_{2},X_{3}\in{\mathbb{R}}^{n\times n}$. Then the Lagrangian associated with ($-{\mathbb{R}}$) given by
@@
Therefore, the dual problem of ($-{\mathbb{R}}$) can be written down as
($-{\mathbb{R}}$’) $\begin{cases}\sup\limits_{X\in{\mathbf{S}}^{2n}}&\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(X_{1}+X_{2})+{\mathscr{A}}_{I}(X_{3}-X_{3}^{\intercal})=b_{R},\\ &{\mathscr{A}}_{R}(X_{3}-X_{3}^{\intercal})-{\mathscr{A}}_{I}(X_{1}+X_{2})=b_{I},\\ &X=\begin{bmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{bmatrix}\succeq 0.\end{cases}$
@@
($-{\mathbb{R}}$) is equivalent to ($-{\mathbb{R}}$) (in the sense that they share the same optimum). As a result, ($-{\mathbb{R}}$) is equivalent to ($-{\mathbb{C}}$). In addition, if $X^{\star}=\left[\begin{smallmatrix}X_{1}^{\star}&(X_{3}^{\star})^{\intercal}\\ X_{3}^{\star}&X_{2}^{\star}\end{smallmatrix}\right]$ is an optimal solution to ($-{\mathbb{R}}$), then $H^{\star}=(X^{\star}_{1}+X_{2}^{\star})+(X^{\star}_{3}-(X_{3}^{\star})^{\intercal})\hbox{\bf{i}}$ is an optimal solution to ($-{\mathbb{C}}$).
@@
Let us denote the optima of ($-{\mathbb{R}}$) and ($-{\mathbb{R}}$) by $v$ and $v^{\prime}$, respectively. Suppose that $Y=\left[\begin{smallmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{smallmatrix}\right]$ is a feasible solution to ($-{\mathbb{R}}$). Then one can easily check that $X\coloneqq\frac{1}{2}Y$ is a feasible solution to ($-{\mathbb{R}}$). Moreover, we have $\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle=\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle$ and it follows $v\leq v^{\prime}$. On the other hand, suppose $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]$ is a feasible solution to ($-{\mathbb{R}}$). We then have
@@
and thus
@@
One can easily see that $Y$ is a feasible solution to ($-{\mathbb{R}}$) and in addition, it holds $\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle=\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle$. Thus $v\geq v^{\prime}$, which proves the equivalence. The latter statement of the theorem is clear from the above arguments.
In contrast to ($-{\mathbb{R}}$), the PSD constraint in ($-{\mathbb{R}}$) is straightforward, and thus no extra affine constraint is required. This is why the conversion ($-{\mathbb{R}}$) is more appealing than ($-{\mathbb{R}}$) from the computational perspective.
@@
A similar reformulation to ($-{\mathbb{R}}$) but for a restricted class of complex SDP relaxations of multiple-input multiple-output detection has appeared in [11].
@@ 3 Application to complex SDP relaxations for CPOPs
In this section, we apply the reformulation ($-{\mathbb{R}}$) to complex SDP relaxations arising from the complex moment-HSOS hierarchy for CPOPs. A CPOP is given by
@@
where $\overline{{\mathbf{z}}}\coloneqq(\overline{z}_{1},\ldots,\overline{z}_{s})$ stands for the conjugate of complex variables ${\mathbf{z}}\coloneqq(z_{1},\ldots,z_{s})$. The functions $f,g_{1},\ldots,g_{t}$ are real-valued polynomials and their coefficients satisfy $b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{b_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$, $g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{g^{i}_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$. The support of $f$ is defined by $\hbox{\rm{supp}}(f)\coloneqq\{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\mid b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$. For $i\in[t]$, $\hbox{\rm{supp}}(g_{i})$ is defined in the same way.
Fix a $d\in{\mathbb{N}}$. Let $y=(y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}})_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}}\subseteq{\mathbb{C}}$ be a sequence indexed by $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}$ and satisfying $y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{y_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$. Let $L_{y}$ be the linear functional defined by
@@
$[{\mathbf{M}}_{d}(y)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}({\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
@@
$[{\mathbf{M}}_{d}(gy)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}(g{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=\sum_{({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})}g_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}y_{{\boldsymbol{\beta}}+{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}+{\boldsymbol{\gamma}}^{\prime}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
Let $d_{0}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ and $d_{i}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ for $i\in[t]$, where ||$\cdot$|| ⋅ | denotes the sum of entries. Let $d_{\min}\coloneqq\max\,\{d_{0},d_{1},\ldots,d_{m}\}$. For any $d\geq d_{\min}$, the $d$-th ($d$ is called the relaxation order) complex moment relaxation for (CPOP) is given by
($-{\mathbb{C}}$) $\begin{cases}\inf\limits_{y}&L_{y}(f)=\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&{\mathbf{M}}_{d}(y)\succeq 0,\\ &{\mathbf{M}}_{d-d_{i}}(g_{i}y)\succeq 0,\quad i\in[t],\\ &y_{\mathbf{0},\mathbf{0}}=1.\end{cases}$
($-{\mathbb{C}}$) and its dual problem form the complex moment-HSOS hierarchy of (CPOP). For more details on this hierarchy, we refer the reader to [10, 15].
@@
(6) $[A^{0}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}1,&\text{if }({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$

### v4

@@ Abstract
This note proposes a new reformulation of complex semidefinite programs (SDP) as real SDPs. As an application, we present an economical reformulation of complex SDP relaxations of complex polynomial optimization problems as real SDPs and derive some further reductions by exploiting inner structure of the complex SDP relaxations. Various numerical examples demonstrate that our new reformulation runs significantly faster than the usual popular reformulation.
@@ 1 Introduction
Complex semidefinite programs (SDP) arise from a diverse set of areas, such as combinatorial optimization [11], optimal power flow [12, 14], quantum information theory [2, 4, 19], and signal processing [13, 16]. In particular, they appear as convex relaxations of complex polynomial optimization problems (CPOP), giving rise to the complex moment-Hermitian-sum-of-squares (moment-HSOS) hierarchy [12, 17, 18]. However, most modern SDP solvers deal with only real SDPs [^1] As far as the author knows, SeDuMi [15], Sdolab [9], and Hypatia [8] are the only solvers that can handle complex SDPs directly.. In order to handle complex SDPs via real SDP solvers, it is mandatory to reformulate complex SDPs as equivalent real SDPs. A popular way [^2] See for instance the online modeling cookbook of the commercial SDP solver MOSEK: https://docs.mosek.com/modeling-cookbook/sdo.html. to do so is to use the equivalent condition
@@
for a Hermitian matrix variable $H=H_{R}+H_{I}\hbox{\bf{i}}\in{\mathbb{C}}^{n\times n}$ with $H_{R}$ and $H_{I}$ being its real and imaginary parts respectively. Note that the right-hand-side constraint in (1) entails certain structure and to feed it to an SDP solver, we need to impose extra affine constraints to the positive semidefinite (PSD) constraint $Y\succeq 0$:
@@
This conversion is quite simple but could be inefficient when $n$ is large. In this note, inspired by Lagrange duality, we propose a new reformulation of complex SDPs as real SDPs. The benefit of this new reformulation is that it contains no extra affine constraints while keeping the same matrix size and hence has lower complexity. In the same manner, we can obtain a new reformulation of complex SDP relaxations of CPOPs as real SDPs. Moreover, by exploiting inner structure of the complex SDP relaxations, we are able to remove a bunch of redundant affine constraints, which leads to an even more economical real reformulation of the complex SDP relaxations. Various numerical experiments (including the complex matrix completion problem, complex SDP relaxations for randomly generated CPOPs, and the alternating current optimal power flow (AC-OPF) problem) confirm our theoretical expectation and demonstrate that the new reformulation is substantially more efficient than the usual popular one. Finally, numerical experiments also indicate that our implementation of the new reformulation with MOSEK [1] runs much faster than the implementation of the original complex formulation with Hypatia [10], probably because MOSEK is commercial and the implementations of SDP algorithms based on real numbers are more mature and robust.
Notation. The symbol $\mathbb{N}$} denotes the set of nonnegative integers. For $n\in{\mathbb{N}}\setminus\{0\}$, let $n]\coloneqq\{1,2,\ldots,n\}$. We write $|A|$ for the cardinality of a set $A$. Let i be the imaginary unit, satisfying $\hbox{\bf{i}}^{2}=-1$. For $d\in{\mathbb{N}}$, let ${\mathbb{N}}^{n}_{d}\coloneqq\{(\alpha_{i})_{i}\in{\mathbb{N}}^{n}\mid\sum_{i=1}^{n}\alpha_{i}\leq d\}$ and let $\omega_{n,d}\coloneqq\binom{n+d}{d}$ be the cardinality of ${\mathbb{N}}^{n}_{d}$. For ${\boldsymbol{\alpha}}=(\alpha_{i})_{i}\in{\mathbb{N}}^{n}_{d}$ and an $n$-tuple of variables ${\mathbf{z}}=\{z_{1},\ldots,z_{n}\}$, let ${\mathbf{z}}^{{\boldsymbol{\alpha}}}\coloneqq z_{1}^{\alpha_{1}}\cdots z_{n}^{\alpha_{n}}$. For a complex number $a$, $\overline{a}$ (resp. ${\mathcal{R}}(a),{\mathcal{I}}(a)$) denotes the conjugate (resp. real part, imaginary part) of $a$, and for a complex vector $v$, $v^{{\mathrm{H}}}$ denotes the conjugate transpose of $v$. For a positive integer $n$, the set of $n\times n$ symmetric (resp. Hermitian) matrices is denoted by ${\mathbf{S}}^{n}$ (resp. ${\mathrm{H}}^{n}$). We use $A\succeq 0$ to indicate that the matrix $A$ is PSD. For $A,B\in{\mathbb{C}}^{n\times n}$, we denote by $\langle A,B\rangle$ the trace inner-product, defined by $\langle A,B\rangle=\hbox{\rm{Tr}}(A^{{\mathrm{H}}}B)$, where $A^{{\mathrm{H}}}$ stands for the conjugate transpose of $A$. For $\in{\mathbb{R}}^{n\times n}, AA^{\intercal}$ stands for the transpose of $A$.
@@
$\langle u,v\rangle_{{\mathbb{R}}}={\mathcal{R}}(u^{{\mathrm{H}}}v)={\mathcal{R}}(u)^{\intercal}{\mathcal{R}}(v)+{\mathcal{I}}(u)^{\intercal}{\mathcal{I}}(v),\quad u,v\in{\mathbb{C}}^{m},$
which will be used in forming the dual complex SDP problem (to obtain a real objective). For $\in{\mathbb{R}}^{m}, u,vuv\langle$ $\rangle\coloneqq u^{\intercal}v$, where $u^{\intercal}$ stands for the transpose of $u$. We associate each $A\in{\mathbb{C}}^{n\times n}$ with a Hermitian matrix ${\mathcal{H}}(A)\coloneqq\frac{1}{2}(A+A^{{\mathrm{H}}})$. One can check that ${\mathcal{R}}(\langle A,H\rangle)=\langle{\mathcal{H}}(A),H\rangle$ for any $H\in{\mathrm{H}}^{n}$.
@@ 2 The real reformulations of complex SDPs
(PSDP-$\mathbb{C}$}) $\begin{cases}\sup\limits_{H\in{\mathrm{H}}^{n}}&\langle C,H\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}(H)=b,\\ &H\succeq 0,\end{cases}$
where $C\in{\mathrm{H}}^{n},b\in{\mathbb{C}}^{m}$. In order to convert (PSDP-$\mathbb{C}$}) to a real SDP, we define two real linear operators ${\mathscr{A}}_{R},{\mathscr{A}}_{I}\colon{\mathbb{R}}^{n\times n}\rightarrow{\mathbb{R}}^{m}$ associated with $\mathscr{A}$} by
@@
respectively. Moreover, assume $H=H_{R}+H_{I}\hbox{\bf{i}},C=C_{R}+C_{I}\hbox{\bf{i}},b=b_{R}+b_{I}\hbox{\bf{i}}$ with $H_{R},H_{I},C_{R},C_{I}\in{\mathbb{R}}^{n\times n}$, $b_{R},b_{I}\in{\mathbb{R}}^{m}$. We can now convert (PSDP-$\mathbb{C}$}) to a real SDP:
(PSDP-$\mathbb{R}$}) $\begin{cases}\sup\limits_{Y\in{\mathbf{S}}^{2n}}&\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(H_{R})+{\mathscr{A}}_{I}(H_{I})=b_{R},\\ &{\mathscr{A}}_{R}(H_{I})-{\mathscr{A}}_{I}(H_{R})=b_{I},\\ &Y=\begin{bmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{bmatrix}\succeq 0.\end{cases}$
As mentioned in the introduction, to feed the PSD constraint in (PSDP-$\mathbb{R}$}) to an SDP solver, we need to include also the extra $n(n+1)$ affine constraints listed in (2), which could be inefficient in practice. Below we show that by taking a dual point of view, we can actually get rid of this issue.
Before formulating the dual problem of (PSDP-$\mathbb{C}$}), we explicitly give the adjoint operator of $\mathscr{A}$}.
@@
Then by the convex duality theory, the dual problem of (PSDP-$\mathbb{C}$}) reads as
(DSDP-$\mathbb{C}$}) $\begin{cases}\inf\limits_{y\in{\mathbb{C}}^{m}}&\langle b,y\rangle_{{\mathbb{R}}}\\ \,\,\,\rm{s.t.}&{\mathscr{A}}^{*}(y)\succeq C.\end{cases}$
@@
Thus, we can convert (DSDP-$\mathbb{C}$}) to a real SDP by using the equivalent condition (1):
(DSDP-$\mathbb{R}$}) $\begin{cases}\inf\limits_{y_{R},y_{I}\in{\mathbb{R}}^{m}}&b_{R}^{\intercal}y_{R}+b_{I}^{\intercal}y_{I}\\ \,\,\,\,\,\,\,\,\rm{s.t.}&\begin{bmatrix}U-C_{R}&-V+C_{I}\\ V-C_{I}&U-C_{R}\end{bmatrix}\succeq 0.\end{cases}$
Let $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]\in{\mathbf{S}}^{2n}$ be the PSD variable of (DSDP-$\mathbb{R}$}) with $X_{1},X_{2},X_{3}\in{\mathbb{R}}^{n\times n}$. Then the Lagrangian associated with (DSDP-$\mathbb{R}$}) is given by
@@
Therefore, the dual problem of (DSDP-$\mathbb{R}$}) can be written down as
(PSDP-$\mathbb{R}$}’) $\begin{cases}\sup\limits_{X\in{\mathbf{S}}^{2n}}&\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle\\ \,\,\,\,\rm{s.t.}&{\mathscr{A}}_{R}(X_{1}+X_{2})+{\mathscr{A}}_{I}(X_{3}-X_{3}^{\intercal})=b_{R},\\ &{\mathscr{A}}_{R}(X_{3}-X_{3}^{\intercal})-{\mathscr{A}}_{I}(X_{1}+X_{2})=b_{I},\\ &X=\begin{bmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{bmatrix}\succeq 0.\end{cases}$
@@
(PSDP-$\mathbb{R}$}) is equivalent to (PSDP-$\mathbb{R}$}) (in the sense that they share the same optimum). As a result, (PSDP-$\mathbb{R}$}) is equivalent to (PSDP-$\mathbb{C}$}). In addition, if $X^{\star}=\left[\begin{smallmatrix}X_{1}^{\star}&(X_{3}^{\star})^{\intercal}\\ X_{3}^{\star}&X_{2}^{\star}\end{smallmatrix}\right]$ is an optimal solution to (PSDP-$\mathbb{R}$}), then $H^{\star}=(X^{\star}_{1}+X_{2}^{\star})+(X^{\star}_{3}-(X_{3}^{\star})^{\intercal})\hbox{\bf{i}}$ is an optimal solution to (PSDP-$\mathbb{C}$}).
@@
Let us denote the optima of (PSDP-$\mathbb{R}$}) and (PSDP-$\mathbb{R}$}) by $v$ and $v^{\prime}$, respectively. Suppose that $Y=\left[\begin{smallmatrix}H_{R}&-H_{I}\\ H_{I}&H_{R}\end{smallmatrix}\right]$ is a feasible solution to (PSDP-$\mathbb{R}$}). Then one can easily check that $X\coloneqq\frac{1}{2}Y$ is a feasible solution to (PSDP-$\mathbb{R}$}). Moreover, we have $\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle=\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle$ and it follows $v\leq v^{\prime}$. On the other hand, suppose $X=\left[\begin{smallmatrix}X_{1}&X_{3}^{\intercal}\\ X_{3}&X_{2}\end{smallmatrix}\right]$ is a feasible solution to (PSDP-$\mathbb{R}$}). Since PSDness is preserved under basis transformation, we have
@@
and since PSDness is preserved under addition, we further have
@@
One can easily see that $Y$ is a feasible solution to (PSDP-$\mathbb{R}$}) and in addition, it holds $\langle C_{R},H_{R}\rangle+\langle C_{I},H_{I}\rangle=\langle C_{R},X_{1}+X_{2}\rangle+\langle C_{I},X_{3}-X_{3}^{\intercal}\rangle$. Thus $v\geq v^{\prime}$, which proves the equivalence. The latter statement of the theorem is clear from the above arguments.
In contrast to (PSDP-$\mathbb{R}$}), the PSD constraint in (PSDP-$\mathbb{R}$}) does not entails any special structure, and thus no extra affine constraint is required to describe it. This is why the conversion (PSDP-$\mathbb{R}$}) is more appealing than (PSDP-$\mathbb{R}$}) from the computational perspective.
@@
A similar reformulation to (PSDP-$\mathbb{R}$}) but for a restricted class of complex SDP relaxations of multiple-input multiple-output detection has appeared in [13].
@@ 3 Application to complex SDP relaxations for CPOPs
In this section, we apply the reformulation (PSDP-$\mathbb{R}$}) to complex SDP relaxations arising from the complex moment-HSOS hierarchy for CPOPs. A CPOP is given by
@@
where $\overline{{\mathbf{z}}}\coloneqq(\overline{z}_{1},\ldots,\overline{z}_{s})$ stands for the conjugate of complex variables ${\mathbf{z}}\coloneqq(z_{1},\ldots,z_{s})$. The functions $f,g_{1},\ldots,g_{t}$ are real-valued polynomials in $\mathbf{z}$} and $\overline{{\mathbf{z}}}$ (their coefficients satisfy $b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{b_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$, $g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{g^{i}_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$). The support of $f$ is defined by $\hbox{\rm{supp}}(f)\coloneqq\{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\mid b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$. For $i\in[t]$, $\hbox{\rm{supp}}(g_{i})$ is defined in the same way.
Fix a $d\in{\mathbb{N}}$. Let $y=(y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}})_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}}$ be a complex sequence indexed by $({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}_{d}^{s}\times{\mathbb{N}}_{d}^{s}$ and satisfying $y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}=\overline{y_{{\boldsymbol{\gamma}},{\boldsymbol{\beta}}}}$. Let $L_{y}$ be the linear functional defined by
@@
${\mathbf{M}}_{d}(y)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}({\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=y_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
@@
${\mathbf{M}}_{d}(gy)]_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\coloneqq L_{y}(g{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}})=\sum_{({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})}g_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}y_{{\boldsymbol{\beta}}+{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}+{\boldsymbol{\gamma}}^{\prime}},\quad\forall{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in{\mathbb{N}}^{s}_{d}.$
Let $d_{0}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:b_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ and $d_{i}\coloneqq\max\,\{|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|:g^{i}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\neq 0\}$ for $i\in[t]$, where $|\cdot|$ denotes the sum of entries. Let $d_{\min}\coloneqq\max\,\{d_{0},d_{1},\ldots,d_{m}\}$. For any $d\geq d_{\min}$, the $d$-th ($d$ is called the relaxation order) complex moment relaxation for (CPOP) is given by
(Mom-$\mathbb{C}$}) $\begin{cases}\inf\limits_{y}&L_{y}(f)=\langle b,y\rangle_{{\mathbb{R}}}\\ \rm{s.t.}&{\mathbf{M}}_{d}(y)\succeq 0,\\ &{\mathbf{M}}_{d-d_{i}}(g_{i}y)\succeq 0,\quad i\in[t],\\ &y_{\mathbf{0},\mathbf{0}}=1.\end{cases}$
(Mom-$\mathbb{C}$}) and its dual problem form the complex moment-HSOS hierarchy of (CPOP). For more details on this hierarchy, we refer the reader to [12, 17].
@@
(6) $A^{0}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}]_{{\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime}}=\begin{cases}1,&\text{if }({\boldsymbol{\beta}}^{\prime},{\boldsymbol{\gamma}}^{\prime})=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),\\ 0,&\text{otherwise}.\end{cases}$